OpenStudy (anonymous):

Solve the equation. 2|7-7x| = 2x+4 A. x=3/2 x=5/6 B. x=3/2 x=5/8 C. x= 5/8 D. x=3/2

OpenStudy (anonymous):

Can you divide both side by 2: 2|7-7x| = 2x+4

OpenStudy (anonymous):

mod(7-7x)=x+2

OpenStudy (anonymous):

when 7-7x=0 x=1

OpenStudy (chihiroasleaf):

|7-7x|=2x+4 7-7x = 2x+4 or 7-7x= -(2x+4) solve these equation

OpenStudy (anonymous):

but u need to check the answers also...because 1st equation is for x less than one...and 2nd one is for x greater than one.

OpenStudy (anonymous):

mod(7-7x)= 7-7x, when x is less than 1. mod(7-7x)=-(7-7x), when x is greater than one.